# Frequency Decomposition-Driven Unsupervised Domain Adaptation for Remote Sensing Image Semantic Segmentation

### v1

@@ 5.9 Model Complexity Analysis
Table 10 shows the computational complexity of all UDA methods discussed. Compared with the baseline Advent, the proposed general backbone GLGAN (Swin-Base) achieved greatly improved segmentation performance with significantly reduced computational complexity as its hierarchical window-based attention has linear computation complexity to input image size (Liu et al., 2021). More specifically, the computational complexity of GLGAN (Swin-Base) is only half that of Advent. Furthermore, GLGAN (Swin-Large) has computational complexity comparable to Advent. However, GLGAN (Swin-Large) requires much more memory to store its parameters. Overall, the proposed GLGAN harvests better performance at the cost of moderately increased computational complexity. The model complexity of FD-GLGAN is between GLGAN (Swin-Base) and GLGAN (Swin-Large), yet it exhibits superior performance. It proved that the HLFD module is of great significance in practical application. Furthermore, although we didn’t conduct FD-GLGAN (Swin-Large) limited by the experimental hardware, its performance holds promise and warrants further exploration.
@@
Table 10: Comparison on computational complexity measured by a $256\times 256$ input on a single NVIDIA GeForce RTX 3090 GPU. Bold values are the best
Model Backbone Complexity (G) Parameters (M) Memory (MB) Speed (FPS)
ResUNet++
ResUNet++
241.82
35.46
4365
25.10
AdasegNet
ResNet101
47.59
42.83
2491
43.31
Advent
ResNet101
47.95
43.16
2491
30.65
GANAI
ResNet101
10.27
44.55
1075
6.30
TriADA
ResNet101
22.15
59.23
1075
4.54
CCAGAN
ResNet101
145.14
104.02
4503
13.04
MBATA-GAN
ResNet101
174.20
143.90
5295
8.11
GLGAN
Swin-Base
26.13
96.72
3639
12.50
Swin-Large
45.97
159.03
4840
9.95
FD-GLGAN
Swin-Base
34.26
138.68
3841
6.22
6 Conclusion
In this work, a novel plug-and-play technology named HLFD has been proposed to learn domain-invariant features through the decomposition of multiscale features from the frequency perspective. To further harness the potential of these decomposed features,a versatile UDA backbone, dubbed GLGAN, has been developed within the GAN framework. GLGAN is adept at extracting both intra- and inter-domain image features, leading to the generation of more precise segmentation maps. Additionally, a novel GLTB-based discriminator called GLDis has been established to enhance the generator’s capability to acquire domain-invariant representations. Extensive experiments on the well-known ISPRS Potsdam and Vaihingen datasets have confirmed that the proposed FD-GLGAN can substantially outperform existing UDA methods while greatly bridging the performance gap between the UDA methods and the supervised methods that utilize labeled data from the target domain. Finally, this work has successfully demonstrated the feasibility of solving the domain shift problem in the UDA tasks using highly efficient GLTB-based backbones and frequency decomposition technology, which will pave the way for future developments of more sophisticated UDA strategies.

### v2

@@ IV-I Model Complexity Analysis
Table XII shows the computational complexity of all UDA methods discussed. Compared with the baseline Advent, the proposed general backbone GLGAN (Swin-Base) achieved greatly improved segmentation performance with significantly reduced computational complexity as its hierarchical window-based attention has linear computation complexity to input image size [40]. More specifically, the computational complexity of GLGAN (Swin-Base) is only half that of Advent. Furthermore, GLGAN (Swin-Large) has computational complexity comparable to Advent. However, GLGAN (Swin-Large) requires much more memory to store its parameters. Overall, the proposed GLGAN harvests better performance at the cost of moderately increased computational complexity. The model complexity of De-GLGAN is between GLGAN (Swin-Base) and GLGAN (Swin-Large), yet it exhibits superior performance. It proved that the HLFD module is of great significance in practical application.
@@
TABLE XII: Comparison on computational complexity measured by a $256\times 256$ input on a single NVIDIA GeForce RTX 3090 GPU. Bold values are the best
Model Backbone Complexity (GFLOPs) Parameters (M) Memory (MB) Speed (FPS)
ResUNet++ ResUNet++ 241.82 35.46 4365 25.10
AdasegNet ResNet101 47.59 42.83 2491 43.31
Advent ResNet101 47.95 43.16 2491 30.65
GANAI ResNet101 10.27 44.55 1075 6.30
TriADA ResNet101 22.15 59.23 1075 4.54
CCAGAN ResNet101 145.14 104.02 4503 13.04
MBATA-GAN ResNet101 174.20 143.90 5295 8.11
MASNet ResNet101 82.34 64.84 3648 18.22
DSSFNet ResNet101 128.52 112.74 4836 11.13
GLGAN Swin-Base 26.13 96.72 3639 12.50
Swin-Large 45.97 159.03 4840 9.95
De-GLGAN Swin-Base 34.26 138.68 3841 6.22
V Conclusion
In this work, a novel decomposition-based UDA approach for remote sensing image semantic segmentation has been proposed. Based on the observation that local detailed features and global contexts are both crucial for reducing domain discrepancy in semantic segmentation, a novel HLFD scheme has been proposed to conduct domain alignment in different subspaces through the decomposition of multiscale features from the frequency perspective. Furthermore, to exploit the global-local cross-domain dependencies, a fully GLGAN has been developed to learn robust domain-invariant representations, leveraging GLTBs. In particular, a novel discriminator, GLDis, has been established to enhance the generator’s representation capability. Finally, a novel UDA framework called De-GLGAN has been established by integrating HLFD and GLGAN for cross-domain semantic segmentation. Extensive experiments on two UDA benchmarks, namely ISPRS Potsdam and Vaihingen, and LoveDA Rural and Urban, have confirmed that the proposed De-GLGAN can substantially outperform existing UDA methods while greatly bridging the performance gap between the existing adversarial training-based UDA approach and the fully-supervised training. Specifically, it can significantly improve the cross-domain transferability and generalization capability of the semantic segmentation model.
In future work, we will investigate the possibility of incorporating the decomposition-based feature learning strategy into fine-tuning visual foundation models on new semantic segmentation datasets. Meanwhile, it is interesting to investigate the integration of the proposed approach with reliable self-training to incorporate category-wise supervision more effectively. Furthermore, it is worth exploring the potential cross-domain applications of GLGAN and the decomposition-based alignment strategy to other tasks, such as image understanding [67, 68] and object detection [69].